A group of 52 people attended a ball game. There were three times as many children as adults in the group. Set up a system of equations that represents the numbers of adults and children who attended the game ans solve the system to find the number of children who were in the group.
Answer by Mathtut(3670)

let the number of adults and children be a and c respectively
:
a+c=52........eq 1
c=3a..........eq 2
:
take c's value from eq 2 and plug it into eq 1
:
a+3a=52
:
4a=52
:
a=13
:
c=13(3)=39